# 2025-04-30

### v1

@@ 1 Introduction
Accurate numerical simulation of physical systems is crucial to predict the system dynamics, carry out control tasks, or undertake design optimization. The high-fidelity solvers based on the numerical discretization of partial differential equations (PDEs), such as the finite volume method or the finite element method, can become prohibitively expensive for high-dimensional problems. This limits their usage, particularly in scenarios requiring real-time predictions. Additionally, if the dynamical system is also parametric in nature, the situation can exacerbate when multiple queries of the high-fidelity solver for varying parameter configurations is required.
To tackle the high computational cost associated with many-query scenarios, there has been significant research and development in designing parametric model reduction strategies [53, 11, 13, 12, 27, 52, 44]. Among them, the reduced basis methods (RBMs) [27, 52, 44] aim at accelerating the computational speed for each parameter query by projecting the underlying governing equations onto a reduced basis space, thereby, requiring to solve only a reduced set of equations. The online query phase is accelerated by introducing a computationally intensive offline phase, where the reduced bases are iteratively adapted along with a sequential collection of high-fidelity solutions corresponding to new parameter locations. The parameter locations at which the high-fidelity solver is queried are typically decided by using a residual-based error estimator such that the accuracy of the resulting ROM improves in an optimal fashion during the offline phase.
@@
In the parametric setting, non-intrusive ROMs are typically built by first querying the high-fidelity solver at a fine set of parameter locations to gather the solution snapshots. Working with such a static set of solution snapshots often requires one to conservatively oversample the parameter space in order to ensure a reasonable accuracy level for the ROM solution. Unfortunately, this inflates the high-fidelity solver queries and greatly burdens the offline phase computations. In literature, limited attention has been given to address this issue. Our work precisely attempts to bridge this gap for non-intrusive parametric model reduction techniques by proposing a novel active learning framework that enables an efficient adaptive sampling of the parameter space during the offline phase.
@@ 1.1 Relation to Prior Work on Active Learning
The existing active learning approaches in the context of data-driven model reduction for parametric systems can be broadly classified into two categories: Error-estimator-based active learning [16, 65, 32], and uncertainity-minimization-based active learning [26, 33]. The error-estimator-based active learning strategies attempt to estimate the non-intrusive ROM error across the parameter domain, thereby, picking new parameter samples in a greedy fashion, at locations corresponding to the highest estimated error values. In order to construct such an error estimator which can be queried at unseen parameter locations, an interpolant is constructed using the ROM error values corresponding to training parameter locations [65, 32]. However, the error behavior across the parameter domain is usually highly irregular, posing a significant challenge to perform accurate interpolation. This was handled in our prior work [32] by a specific design of the error estimator and by performing interpolation using a kernel-based shallow neural network (KSNN) with Matérn class of kernels, trained via an alternating dual-staged iterative training procedure for improved accuracy. The repeated construction of the error estimator at each iteration of the active learning procedure results in an additional computational overhead in the offline phase. Moreover, without restriction to a certain class of model reduction methods [32], it is in general necessary to iteratively construct the ROM at every active learning iteration, which could further result in additional computational efforts.
@@
There are also other noticeable works on active learning [1, 14] and reduction of parameter spaces [18, 17, 21]. However, they do not focus on developing their ideas and applying them to the general setting of data-driven model reduction for parametric systems. Authors in [1] provide a Christoffel adaptive sampling strategy for function approximation using deep neural networks, which can be used to approximate smooth, multivariate, output quantities of interest from parametrized PDEs by using limited data samples. Authors in [14] propose neural Galerkin schemes with active learning that allows sequential updates of a neural network parameters following the Dirac-Frenkel variational principle with an adaptive data collection strategy in time for efficient time-integration of PDEs. The active subspaces approach [18, 17] aim at identifying a set of important directions in the parameter/input domain by performing sensitivity analysis of any function of interest, thereby, reducing the parameter space dimensionality. Authors in [21] have used the active subspaces approach to improve the accuracy of POD modal coefficients’ approximation with ridge functions. However, there is no effort to dynamically grow the training snapshot set by identification of critical locations in the parameter domain, and no notion of active learning, which is precisely the focus of our work.
@@ 1.2 Our Contributions
In contrast to existing approaches, we propose a new paradigm to undertake active learning, wherein, important parameter locations are detected using the distance between solution subspaces. More precisely, our active learning framework relies on measuring the similarity between parameter-specific solution fields by evaluating the distance between their corresponding POD subspaces that encompasses the respective solution behavior. The active learning iterations progressively reduce the gap between parameter-specific solution subspaces. As a result, we observe that the ROM accuracy improves. It is non-trivial to compare the distance between subspaces with different dimensions. We provide a detailed discussion regarding this and also propose a new distance measure and show that it is a metric. Moreover, this distance measure can be repeatedly evaluated for numerous subspace pairs in an computationally efficient fashion, which makes it possible to be efficiently used during the active learning setup. To the best of our knowledge, such an active learning approach, referred to as the subspace-distance-enabled active learning (SDE-AL), has not been previously investigated.
The proposed SDE-AL framework can be deployed in two ways: (a) with a user-defined maximum feasible computational budget for parametric queries of the full-order model (FOM) solver, or (b) with user-defined tolerance levels for the maximal subspace distance between the selected parameter locations and the maximal ROM error. For comparison with the user-specified error tolerance, the underlying ROM error can be evaluated by constructing an error estimator that provides error predictions at unseen parameter locations. In this variant of the algorithm, the error estimator is only used as a stopping criterion in the final stage of the active learning iterations. Moreover, unlike existing error-estimator-based active learning approaches, the central strategy for picking new parameter locations via SDE-AL is independent of the error estimator employed. Furthermore, the SDE-AL approach is general-purpose, meaning it can be flexibly used in conjugation with any non-intrusive ROM. In this article, we showcase the usability of the SDE-AL strategy for POD-KSNN [32] and POD-NN [28, 59] ROMs. Their active-learning-driven counterparts are respectively referred to as SDE-ActLearn-POD-KSNN and SDE-ActLearn-POD-NN ROMs.
@@ 2 Measures of Similarity Between Linear Subspaces
Consider two linear subspaces $\mathcal{X}$ and $\mathcal{Y}$ with their respective orthonormal bases given by the columns of matrices $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n\times q}$. Consider the SVD of the matrix with inner products between the two sets of orthonormal bases, i.e., $X^{\top}Y=W_{l}\Sigma_{XY}W_{r}^{\top}$. Here, $W_{l}$ and $W_{r}$ are orthogonal matrices and $\Sigma_{XY}$ is a $p\times q$ diagonal matrix with singular values $\sigma^{XY}_{1},\dots,\sigma^{XY}_{s}$ in decreasing order and $s=\min(p,q)$. The cosines of the principal angles between the subspaces are given by [58, 64],
@@
To ease further discussions and assist the reader, we formally define what we mean by a metric, premetric, vector norm, and matrix norm. For more details, refer the exposition in [54, 30, 56].
@@
Let $S$ be a set. A metric on $S$ is a function $\rho$ of two arguments, defined on $S$, $s_{1},s_{2},s_{3}\in S$, which satisfies the following conditions:
@@
Let $S$ be a set. A premetric on $S$ is a function $\rho$ of two arguments, defined on $S$, $s_{1},s_{2}\in S$, which satisfies the following conditions:
@@
Let $V$ be a vector space over $\mathbb{R}$. A function $\|\cdot\|:V\rightarrow\mathbb{R}$ is a vector norm if for all $v_{1},v_{2}\in V$ the following axioms are satisfied:
@@
If $\mathcal{X}$ and $\mathcal{Y}$ are linear subspaces with their respective orthonormal bases given by the columns of matrices $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n\times q}$, then the distance measure $\mathcal{D}_{1}(\mathcal{X},\mathcal{Y})$ from eq. 2.2 is only a premetric for $p\neq q$.
@@
$\mathcal{D}_{1}(\mathcal{X},\mathcal{Y})=0\iff\mathcal{X}\subseteq\mathcal{Y}\text{ or }\mathcal{Y}\subseteq\mathcal{X}.$ (2.3)
@@
$\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{X}_{2})\leq\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{Y}_{1})+\mathcal{D}_{1}(\mathcal{X}_{2},\mathcal{Y}_{1}).$ (2.4)
Due to $\mathcal{X}_{1},\mathcal{X}_{2}\subset\mathcal{Y}_{1}$, we know that $\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{Y}_{1})=\mathcal{D}_{1}(\mathcal{X}_{2},\mathcal{Y}_{1})=0$. Following the triangle inequality, this renders $\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{X}_{2})=0$, which is a contradiction. This proves that $\mathcal{D}_{1}(\cdot,\cdot)$ does not satisfy the triangle inequality for $p\neq q$, and hence it is not a metric, but only a premetric as shown before. Note that in this example, we considered linear subspaces of dimensions $1$ and $2$, resulting in distance evaluations between subspaces with $p=1$, $q=2$, as well as with $p=q=1$. Please refer Remark 2.1 for the scenario where all distance evaluations between linear subspaces are restricted to subspaces with $p=q$. ∎
@@
$\|X^{\top}Y\|_{F}=\sqrt{\sum_{i=1}^{p}\sum_{j=1}^{q}|\mathbf{x}_{i}^{\top}\mathbf{y}_{j}|^{2}}=\sqrt{\operatorname{tr}((X^{\top}Y)^{\top}(X^{\top}Y))}.$ (2.5)
@@
$\|X^{\top}Y\|_{F}$ $=\sqrt{\operatorname{tr}(W_{l}\Sigma_{XY}W_{r}^{\top})^{\top}(W_{l}\Sigma_{XY}W_{r}^{\top}))}=\sqrt{\operatorname{tr}(W_{r}\Sigma_{XY}^{\top}\Sigma_{XY}W_{r}^{\top})}=\sqrt{\sum_{k=1}^{s}(\sigma^{XY}_{k})^{2}}.$ (2.6)
@@
$\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]$ $=\operatorname{tr}[(XX^{\top}-YY^{\top})^{2}]$ (2.10)
$=\operatorname{tr}(XX^{\top}+YY^{\top}-XX^{\top}YY^{\top}-YY^{\top}XX^{\top})$ (2.11)
$=\operatorname{tr}(XX^{\top})+\operatorname{tr}(YY^{\top})-2\operatorname{tr}(XX^{\top}YY^{\top})$ (2.12)
$=p+q-2\operatorname{tr}(Y^{\top}XX^{\top}Y)$ (2.13)
$=p+q-2\|X^{\top}Y\|_{F}^{2}$ (2.14)
$=p+q-2\tilde{\mathcal{D}}^{2}(\mathcal{X},\mathcal{Y}).$ (2.15)
$\implies\tilde{\mathcal{D}}(\mathcal{X},\mathcal{Y})$ $=\sqrt{\frac{1}{2}\Big{(}p+q-\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]\Big{)}}.$ (2.16)
@@
If $\mathcal{X}$ and $\mathcal{Y}$ are linear subspaces with their respective orthonormal bases given by the columns of matrices $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n\times q}$, then the distance measure $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ from eq. 2.17 is a metric for all $\0p,q\in\mathbb{N}\backslash 0$.
@@
Let us systematically ensure all the properties of a metric. Consider a scenario when the subspaces are identical, i.e., $\mathcal{X}=\mathcal{Y}$. This means that $p=q$ and the bases matrices $X=Y$. Due to the orthonormalized bases, $\tilde{\mathcal{D}}(\mathcal{X},\mathcal{X})=\sqrt{p}=\sqrt{q}$, which means $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=0$. Hence, this would ensure the identity of indiscernibles, if we can also prove that $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=0$ holds only when the subspaces are identical. We defer this part for later. Next, it is easy to see that the distance measure is symmetric,
$\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=\sqrt{\max(p,q)-\|X^{\top}Y\|^{2}_{F}}=\sqrt{\max(q,p)-\|Y^{\top}X\|^{2}_{F}}=\mathcal{D}_{2}(\mathcal{Y},\mathcal{X})$ (2.18)
@@
$\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Y})=\max(p,q)-\tilde{\mathcal{D}}^{2}(\mathcal{X},\mathcal{Y}).$ (2.19)
@@
$\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Y})$ $=\max(p,q)-\frac{1}{2}\Big{(}p+q-\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]\Big{)}.$ (2.20)
@@
Since the right-hand side of eq. 2.21 is non-negative, it allows us to see that $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})\geq 0$. Moreover, from eq. 2.21 it is also clear that $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=0$, only when $p=q$ and the projectors $P_{\mathcal{X}}=P_{\mathcal{Y}}$, which is only possible when $\mathcal{X}=\mathcal{Y}$. So far, we have successfully shown that properties $1-3$ from Definition 2.1 hold.
Let $\mathbf{v}_{p},\mathbf{v}_{q}\in\mathbb{R}^{n}$ such that the first $p$ and $q$ elements for $\mathbf{v}_{p}$ and $\mathbf{v}_{q}$ respectively are ones and all others are zeros. This allows us to write the following,
$\|\mathbf{v}_{p}-\mathbf{v}_{q}\|_{1}=|p-q|.\,$ (2.22)
@@
$\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]$ $=\operatorname{tr}[(XX^{\top}-YY^{\top})^{2}]$ (2.24)
$=\operatorname{tr}[(XX^{\top}-YY^{\top})(XX^{\top}-YY^{\top})]$ (2.25)
$=\operatorname{tr}[(XX^{\top}-YY^{\top})^{\top}(XX^{\top}-YY^{\top})]$ (2.26)
$=\|XX^{\top}-YY^{\top}\|_{F}.$ (2.27)
@@
where the columns of matrix $Z\in\mathbb{R}^{n\times r}$ correspond to the orthonormal bases spanning a linear subspace $\mathcal{Z}$.
Following eqs. 2.21 and 2.28, the squared distance between the subspaces $\mathcal{X},\mathcal{Y},\mathcal{Z}$ can be written as,
@@
where $\mathcal{D}_{2}(\mathcal{X},\mathcal{Z})$ and $\mathcal{D}_{2}(\mathcal{Z},\mathcal{Y})$ give the distance between subspaces $\mathcal{X},\mathcal{Z}$ and $\mathcal{Z},\mathcal{Y}$ respectively. By adding together eqs. 2.23 and 2.32 we obtain,
$|p-q|+\|P_{\mathcal{X}}-P_{\mathcal{Y}}\|_{F}\leq|p-r|+\|P_{\mathcal{X}}-P_{\mathcal{Z}}\|_{F}+|r-q|+\|P_{\mathcal{Z}}-P_{\mathcal{Y}}\|_{F}.$ (2.36)
Using eqs. 2.33, 2.34 and 2.35, this takes the following form,
$2\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Y})$ $\leq 2\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Z})+2\mathcal{D}_{2}^{2}(\mathcal{Z},\mathcal{Y}).$ (2.37)
$\implies\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ $\leq\sqrt{\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Z})+\mathcal{D}_{2}^{2}(\mathcal{Z},\mathcal{Y})}.$ (2.38)
$\implies\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ $\leq\mathcal{D}_{2}(\mathcal{X},\mathcal{Z})+\mathcal{D}_{2}(\mathcal{Z},\mathcal{Y}).$ (2.39)

### v2

# Subspace-Distance-Enabled Active Learning for Efficient Data-Driven Model Reduction of Parametric Dynamical Systems

@@ 1 Introduction
Accurate numerical simulation of physical systems is crucial to predict the system dynamics, carry out control tasks, or undertake design optimization. The high-fidelity solvers based on the numerical discretization of partial differential equations (PDEs), such as the finite volume method or the finite element method, can become prohibitively expensive for high-dimensional problems. This limits their usage, particularly in scenarios requiring real-time predictions. Additionally, if the dynamical system is also parametric in nature, the situation can exacerbate when multiple queries of the high-fidelity solver for varying parameter configurations are required.
To tackle the high computational cost associated with many-query scenarios, there has been significant research and development in designing parametric model reduction strategies [53, 11, 13, 12, 27, 52, 44]. Among them, the reduced basis methods (RBMs) [27, 52, 44] aim at accelerating the computational speed for each parameter query by projecting the underlying governing equations onto a reduced basis space, thereby, requiring to solve only a reduced set of equations. The online query phase is accelerated by introducing a computationally intensive offline phase, where the reduced bases are iteratively adapted along with a sequential collection of high-fidelity solutions corresponding to important parameter locations. The parameter locations at which the high-fidelity solver is queried are typically decided by using a residual-based error estimator such that the accuracy of the resulting ROM improves in an optimal fashion during the offline phase.
@@
In the parametric setting, non-intrusive ROMs are typically built by first querying the high-fidelity solver at a fine set of parameter locations to gather the solution snapshots. Working with such a static set of solution snapshots often requires one to conservatively oversample the parameter space in order to ensure a reasonable accuracy level for the ROM solution. Unfortunately, this inflates the high-fidelity solver queries and greatly burdens the offline phase computations. In literature, limited attention has been paid to addressing this issue. Our work precisely attempts to bridge this gap for non-intrusive parametric model reduction techniques by proposing a novel active learning framework that enables an efficient adaptive sampling of the parameter space during the offline phase.
@@ 1.1 Relation to Prior Work on Active Learning
The existing active learning approaches in the context of data-driven model reduction for parametric systems can be broadly classified into two categories: Error-estimator-based active learning [16, 65, 32], and uncertainity-minimization-based active learning [26, 33]. The error-estimator-based active learning strategies attempt to estimate the non-intrusive ROM error across the parameter domain, thereby, picking new parameter samples in a greedy fashion, at locations corresponding to the highest estimated error values. In order to construct such an error estimator which can be queried at unseen parameter locations, an interpolant is constructed using the ROM error values corresponding to training parameter locations [65, 32]. However, the error behavior across the parameter domain is usually highly irregular, posing a significant challenge to perform accurate interpolation. This was handled in our prior work [32] by a specific design of the error estimator and by performing interpolation using a kernel-based shallow neural network (KSNN) with the Matérn class of kernels, trained via an alternating dual-staged iterative training procedure for improved accuracy. The repeated construction of the error estimator at each iteration of the active learning procedure results in an additional computational overhead in the offline phase. Moreover, without restriction to a certain class of model reduction methods [32], it is in general necessary to iteratively construct the ROM at every active learning iteration, which could further result in additional computational efforts.
@@
There are also other noticeable works on active learning [1, 14] and reduction of parameter spaces [18, 17, 21]. However, they do not focus on developing their ideas and applying them to the general setting of data-driven model reduction for parametric systems. Authors in [1] provide a Christoffel adaptive sampling strategy for function approximation using deep neural networks, which can be used to approximate smooth, multivariate, output quantities of interest from parametrized PDEs by using limited data samples. Authors in [14] propose neural Galerkin schemes with active learning that allows sequential updates of a neural network parameters following the Dirac-Frenkel variational principle with an adaptive data collection strategy in time for efficient time-integration of PDEs. The active subspaces approach [18, 17] aims at identifying a set of important directions in the parameter/input domain by performing sensitivity analysis of any function of interest, thereby, reducing the parameter space dimensionality. Authors in [21] have used the active subspaces approach to improve the accuracy of POD modal coefficients’ approximation with ridge functions. However, there is no effort to dynamically grow the training snapshot set by identification of critical locations in the parameter domain, and no notion of active learning, which is precisely the focus of our work.
@@ 1.2 Our Contributions
In contrast to existing approaches, we propose a new paradigm to undertake active learning, wherein important parameter locations are detected using the distance between solution subspaces. More precisely, our active learning framework relies on measuring the similarity between parameter-specific solution fields by evaluating the distance between their corresponding POD subspaces that encompasses the respective solution behavior. The active learning iterations progressively reduce the gap between parameter-specific solution subspaces. As a result, we observe that the ROM accuracy improves. It is non-trivial to compare the distance between subspaces with different dimensions. We provide a detailed discussion regarding this and also propose a new distance measure and show that it is a metric. Moreover, this distance measure can be repeatedly evaluated for numerous subspace pairs in an computationally efficient fashion, which makes it possible to be efficiently used during the active learning setup. To the best of our knowledge, such an active learning approach, referred to as the subspace-distance-enabled active learning (SDE-AL), has not been previously investigated.
The proposed SDE-AL framework can be deployed in two ways: (a) with a user-defined maximum feasible computational budget for parametric queries of the full-order model (FOM) solver, or (b) with user-defined tolerance levels for the maximal subspace distance between the selected parameter locations and the maximal ROM error. For comparison with the user-specified error tolerance, the underlying ROM error can be evaluated by constructing an error estimator that provides error predictions at unseen parameter locations. In this variant of the algorithm, the error estimator is only used as a stopping criterion in the final stage of the active learning iterations. Moreover, unlike existing error-estimator-based active learning approaches, the central strategy for picking new parameter locations via SDE-AL is independent of the error estimator employed. Furthermore, the SDE-AL approach is general-purpose, meaning it can be flexibly used in conjugation with any non-intrusive reduced-order modeling method. In this article, we showcase the usability of the SDE-AL strategy for POD-KSNN [32] and POD-NN [28, 59] ROMs. Their active-learning-driven counterparts are respectively referred to as SDE-ActLearn-POD-KSNN and SDE-ActLearn-POD-NN ROMs.
@@ 2 Measures of Similarity Between Linear Subspaces
Consider two linear subspaces $\mathcal{X}$ and $\mathcal{Y}$ of $\mathbb{R}^{n}$ with their respective orthonormal bases given by the columns of matrices $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n\times q}$. Consider the SVD of the matrix with inner products between the two sets of orthonormal bases, i.e., $X^{\top}Y=W_{l}\Sigma_{XY}W_{r}^{\top}$. Here, $W_{l}$ and $W_{r}$ are orthogonal matrices and $\Sigma_{XY}$ is a $p\times q$ diagonal matrix with singular values $\sigma^{XY}_{1},\dots,\sigma^{XY}_{s}$ in decreasing order and $s=\min(p,q)$. The cosines of the principal angles between the subspaces are given by [58, 64],
@@
To ease further discussions and assist the reader, we formally define what we mean by a metric, premetric, vector norm, and matrix norm. For more details, refer to the exposition in [54, 30, 56].
@@
Let $S$ be a set. A metric on $S$ is a function $\rho$ of two arguments, defined on $S$, where $s_{1},s_{2},s_{3}\in S$, which satisfies the following conditions:
@@
Let $S$ be a set. A premetric on $S$ is a function $\rho$ of two arguments $s_{1},s_{2}\in S$, which satisfies the following conditions:
@@
Let $V$ be a vector space over $\mathbb{R}$. A function $\|\cdot\|:V\rightarrow\mathbb{R}$ is a vector norm if for all $v_{1},v_{2}\in V$, the following axioms are satisfied:
@@
If $\mathcal{X}$ and $\mathcal{Y}$ are linear subspaces with their respective orthonormal bases given by the columns of matrices $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n\times q}$, then the distance measure $\mathcal{D}_{1}(\mathcal{X},\mathcal{Y})$ from eq. 2.2 is a premetric for $p\neq q$, but in general not a metric.
@@
$\mathcal{D}_{1}(\mathcal{X},\mathcal{Y})=0\iff\mathcal{X}\subseteq\mathcal{Y}\text{ or }\mathcal{Y}\subseteq\mathcal{X}.$
@@
$\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{X}_{2})\leq\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{Y}_{1})+\mathcal{D}_{1}(\mathcal{X}_{2},\mathcal{Y}_{1}).$
Due to $\mathcal{X}_{1},\mathcal{X}_{2}\subset\mathcal{Y}_{1}$, we know that $\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{Y}_{1})=\mathcal{D}_{1}(\mathcal{X}_{2},\mathcal{Y}_{1})=0$. Following the triangle inequality, this renders $\mathcal{D}_{1}(\mathcal{X}_{1},\mathcal{X}_{2})=0$, which is a contradiction. This proves that $\mathcal{D}_{1}(\cdot,\cdot)$ does not satisfy the triangle inequality for $p\neq q$, and hence it is not a metric, but only a premetric as shown before. ∎
Note that for the example in the proof above, we considered linear subspaces of dimensions $1$ and $2$, resulting in distance evaluations between subspaces with $p=1$, $q=2$, as well as with $p=q=1$. Please refer Remark 2.1 for the scenario where all distance evaluations between linear subspaces are restricted to subspaces with $p=q$.
@@
$\|X^{\top}Y\|_{F}=\sqrt{\sum_{i=1}^{p}\sum_{j=1}^{q}|\mathbf{x}_{i}^{\top}\mathbf{y}_{j}|^{2}}=\sqrt{\operatorname{tr}((X^{\top}Y)^{\top}(X^{\top}Y))}.$
@@
$\|X^{\top}Y\|_{F}$ $=\sqrt{\operatorname{tr}(W_{l}\Sigma_{XY}W_{r}^{\top})^{\top}(W_{l}\Sigma_{XY}W_{r}^{\top}))}=\sqrt{\operatorname{tr}(W_{r}\Sigma_{XY}^{\top}\Sigma_{XY}W_{r}^{\top})}=\sqrt{\sum_{k=1}^{s}(\sigma^{XY}_{k})^{2}}.$
@@
$\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]$ $=\operatorname{tr}[(XX^{\top}-YY^{\top})^{2}]$
$=\operatorname{tr}(XX^{\top}+YY^{\top}-XX^{\top}YY^{\top}-YY^{\top}XX^{\top})$
$=\operatorname{tr}(XX^{\top})+\operatorname{tr}(YY^{\top})-2\operatorname{tr}(XX^{\top}YY^{\top})$
$=p+q-2\operatorname{tr}(Y^{\top}XX^{\top}Y)$
$=p+q-2\|X^{\top}Y\|_{F}^{2}$
$=p+q-2\tilde{\mathcal{D}}^{2}(\mathcal{X},\mathcal{Y}).$
$\implies\tilde{\mathcal{D}}(\mathcal{X},\mathcal{Y})$ $=\sqrt{\frac{1}{2}\Big{(}p+q-\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]\Big{)}}.$
@@
If $\mathcal{X}$ and $\mathcal{Y}$ are linear subspaces with their respective orthonormal bases given by the columns of matrices $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n\times q}$, then the distance measure $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ from eq. 2.6 is a metric for all $\0p,q\in\mathbb{N}\backslash\{0\}$.
@@
Let us systematically ensure all the properties of a metric. Consider a scenario when the subspaces are identical, i.e., $\mathcal{X}=\mathcal{Y}$. This means that $p=q$ and $\operatorname*{range}(X)=\operatorname*{range}(Y)$. Due to the orthonormalized bases, $\tilde{\mathcal{D}}(\mathcal{X},\mathcal{X})=\sqrt{p}=\sqrt{q}$, which means $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=0$. Hence, this would ensure the identity of indiscernibles, if we can also prove that $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=0$ holds only when the subspaces are identical. We defer this part for later. Next, it is easy to see that the distance measure is symmetric,
$\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=\sqrt{\max(p,q)-\|X^{\top}Y\|^{2}_{F}}=\sqrt{\max(q,p)-\|Y^{\top}X\|^{2}_{F}}=\mathcal{D}_{2}(\mathcal{Y},\mathcal{X}).$
Moreover, $\mathcal{D}_{2}$ is non-negative by construction, i.e., $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})\geq 0$.
@@
$\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Y})=\max(p,q)-\tilde{\mathcal{D}}^{2}(\mathcal{X},\mathcal{Y}).$
@@
$\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Y})$ $=\max(p,q)-\frac{1}{2}\Big{(}p+q-\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]\Big{)}.$
@@
It is clear from eq. 2.7 that $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})=0$, only when $p=q$ and the projectors $P_{\mathcal{X}}=P_{\mathcal{Y}}$, which is only possible when $\mathcal{X}=\mathcal{Y}$. So far, we have successfully shown that properties $1-3$ from Definition 2.1 hold.
Let $\mathbf{v}_{p},\mathbf{v}_{q}\in\mathbb{R}^{n}$ be such that the first $p$ and $q$ elements for $\mathbf{v}_{p}$ and $\mathbf{v}_{q}$ respectively are ones and all others are zeros. This allows us to write the following,
$\|\mathbf{v}_{p}-\mathbf{v}_{q}\|_{1}=|p-q|.\,$
@@
$\operatorname{tr}[(P_{\mathcal{X}}-P_{\mathcal{Y}})^{2}]$ $=\operatorname{tr}[(XX^{\top}-YY^{\top})^{2}]$
$=\operatorname{tr}[(XX^{\top}-YY^{\top})(XX^{\top}-YY^{\top})]$
$=\operatorname{tr}[(XX^{\top}-YY^{\top})^{\top}(XX^{\top}-YY^{\top})]$
$=\|XX^{\top}-YY^{\top}\|_{F}.$
@@
where the columns of matrix $Z\in\mathbb{R}^{n\times r}$ correspond to the orthonormal basis spanning a linear subspace $\mathcal{Z}$.
Following eqs. 2.7 and 2.9, the squared distance between the subspaces $\mathcal{X},\mathcal{Y},\mathcal{Z}$ can be written as,
@@
where $\mathcal{D}_{2}(\mathcal{X},\mathcal{Z})$ and $\mathcal{D}_{2}(\mathcal{Z},\mathcal{Y})$ give the distance between subspaces $\mathcal{X},\mathcal{Z}$ and $\mathcal{Z},\mathcal{Y}$ respectively. By adding eqs. 2.8 and 2.10, we obtain
$|p-q|+\|P_{\mathcal{X}}-P_{\mathcal{Y}}\|_{F}\leq|p-r|+\|P_{\mathcal{X}}-P_{\mathcal{Z}}\|_{F}+|r-q|+\|P_{\mathcal{Z}}-P_{\mathcal{Y}}\|_{F}.$
Using eqs. 2.11, 2.12 and 2.13, this takes the following form,
$2\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Y})$ $\leq 2\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Z})+2\mathcal{D}_{2}^{2}(\mathcal{Z},\mathcal{Y}).$
$\implies\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ $\leq\sqrt{\mathcal{D}_{2}^{2}(\mathcal{X},\mathcal{Z})+\mathcal{D}_{2}^{2}(\mathcal{Z},\mathcal{Y})}.$
$\implies\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ $\leq\mathcal{D}_{2}(\mathcal{X},\mathcal{Z})+\mathcal{D}_{2}(\mathcal{Z},\mathcal{Y}).$